OpenStudy (anonymous):

write an equation of the sine function with amplitude 5 period and phase shift –pi

OpenStudy (anonymous):

y=2sin(x-pi)-1 y=.5sin8(x)+3

OpenStudy (anonymous):

thats none of my choices

OpenStudy (anonymous):

ok give me your choies

OpenStudy (anonymous):

y= tan(4x-pi/4)+1 y= tan (4x-4pi)+1 y= tan (x/4-pi)+1 y= tan (4x+4pi)+1

OpenStudy (anonymous):

i was guessing the first one...

OpenStudy (anonymous):

wait wrong question...

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

those choices are for the question; write an equation of the tangent function with period pi/4, phase shift pi,and vertical shift 1

OpenStudy (anonymous):

idk about these just trust your instinc

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob